OpenStudy (anonymous):

How Do I Solve " -2x^2=4x-3 " ?

OpenStudy (saifoo.khan):

If u want to get less confuse, use the quadratic formua;

OpenStudy (saifoo.khan):

Formula.

OpenStudy (anonymous):

Don't Remember It .

OpenStudy (dumbcow):

set it equal to zero first...................then plug in coefficients into following quadratic formula \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

So How Would I Set It Up ?

OpenStudy (dumbcow):

identify your coefficients, a is with x^2, b is with x, c is the constant

OpenStudy (dumbcow):

make sense? ...............whats a,b,c for this equation?

OpenStudy (anonymous):

u would have to rearrange for it to become this way \[ax^2+bx+c=0\]

OpenStudy (anonymous):

Can You Show Me How Too Set It Up With The Numbers In Place ? And I'll Try Too Solve It MySelf .

OpenStudy (anonymous):

compare -2x^2=4x-3 with ax^2+bx+c=0 rearrange it to get -2x^2-4x+3=0 and now let us do some substitution a=-2 b=-4 c=3 i hope is now understood